ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 91, 1995

Abels, H., G. A. Margulis and G. A. Soifer Semigroups containing proximal linear maps

1

Aljadeff, E. and D. J. S. Robinson Solution of the twisting problem for skew group algebras

409

Barcelo, H. and A. Goupil Non-broken circuits of reflection groups and factorization in D_{n}

285

Berele, A. and A. Regev On the codimensions of the verbally prime P.I. algebras

239

Blair, D. E., T. Koufogiorgos and B. J. Papantoniou Contact metric manifolds satisfying a nullity condition

189

Blasco, J. L. and C. Ivorra Injective spaces of real-valued functions with the Baire property

341

Brehm, U. and G. Schild Realizability of the torus and the projective plane in ℝ^{4}

249

Chen, B. Y. Submanifolds in de Sitter space-time satisfying ΔH =

λH 373

Chuaqui, M. and B. Osgood An extension of a theorem of Gehring and Pommerenke

393

Efrat, I. A Galois-theoretic characterization of p-adically closed fields

273

Gol’dshtein, V., L. Gurov and A. Romanov Homeomorphisms that induce monomorphisms of Sobolev spaces

31

Goupil, A. See H. Barcelo

Gurov, L. See V. Gol’dshtein

Handel, M., B. Kitchens and D. J. Rudolph Metrics and entropy for non-compact spaces

253

Host, B. Nombres normaux, entropie, translations  419

Ivorra, C. See J. L. Blasco

Kitchens, B. See M. Handel

Knörr, R., W. Lempken and B. Thielcke The S_{3}-conjecture for solvable groups

61

Koufogiorgos, T. See D. E. Blair

Lempken, W. See R. Knörr Ma, W. and D. Minda Behavior of domain constants under

conformal mappings

157

Margulis, G. A. See H. Abels Mason, A. W. Quotients of the congruence kernels of SL_{2} over

arithmetic Dedekind domains

77

Mauldin, R. D. and M. G. Monticino Randomly generated distributions

215

Minda, D. See W. Ma Monticino, M. G. See R. D. Mauldin Osgood, B. See M. Chuaqui Papantoniou, B. J. See D. E. Blair Pollicott, M. One-dimensional maps via complex analysis in several

variables

317

Rawat, R. and A. Sitaram The injectivity of the Pompeiu transform and L^{p}-analogues of the Wiener–Tauberian theorem

307

Regev, A. See A. Berele Robinson, D. J. S. See E. Aljadeff Romanov, A. See V. Gol’dshtein Rudolph, D. J. See M. Handel Schild, G. See U. Brehm Schmutz, E. The order of a typical matrix with entries in a finite

field

349

Sitaram, A. See R. Rawat Soifer, G. A. See H. Abels Thaler, M. A limit theorem for the Perron–Frobenius operator of

transformations on [0, 1] with indifferent fixed points

111

Thielcke, B. See R. Knörr

Trow, P. Decompositions of finite-to-one factor maps  129
Tylli, H.-O. The essential norm of an operator is not self-dual  93

Xu, Q. Some results related to interpolation on Hardy spaces of regular martingales 173